# Soliton Amplification in the Korteweg-de Vries Equation by Multiplicative Forcing

### v1

@@ 4 Short-time control
[…]
@@
Corollary 4.4 implies that
@@
Using furthermore that $\int_{0}^{\delta}(\delta-s)^{-1/2}{\mathrm{d}}s=2\sqrt{\delta}$, we find that
@@
$P_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)=\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t_{\diamond})}^{1}\big{\rangle}\partial_{x}\phi_{c(t_{\diamond})}+\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t)}^{2}\big{\rangle}\partial_{c}\phi_{c(t_{\diamond})}$
where $\mathcal{L}_{c(t_{\diamond})}\partial_{x}\phi_{c(t_{\diamond})}=0$ and $\mathcal{L}_{c(t_{\diamond})}\partial_{c}\phi_{c(t_{\diamond})}=\partial_{x}\phi_{c(t_{\diamond})}$. Thus
@@
$+\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t_{\diamond})}^{2}\big{\rangle}(\partial_{c}\phi_{c(t_{\diamond})}+(\delta-s)\partial_{x}\phi_{c(t_{\diamond})}),$
@@
$\|e^{(r-s)\mathcal{L}_{c(t_{\diamond})}}P_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{H^{1}_{\overline{\mathbf{w}}}}\leq$ $\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}\|\eta_{c(t_{\diamond})}^{1}\|_{L^{2}_{-{\overline{\mathbf{w}}}}}\|\partial_{x}\phi_{c(t_{\diamond})}\|_{H^{1}_{\overline{\mathbf{w}}}}$
$+\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}\|\eta_{c(t_{\diamond})}^{2}\|_{L^{2}_{-{\overline{\mathbf{w}}}}}\|\partial_{c}\phi_{c(t_{\diamond})}\|_{H^{1}_{\overline{\mathbf{w}}}}$
$+\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}\|\eta_{c(t_{\diamond})}^{2}\|_{L^{2}_{-{\overline{\mathbf{w}}}}}(r-s)\|\partial_{x}\phi_{c(t_{\diamond})}\|_{H^{1}_{\overline{\mathbf{w}}}}$
@@
$\left\|\int_{0}^{r}e^{(r-s)\mathcal{L}_{c(t_{\diamond})}}P_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)\ {\mathrm{d}}s\right\|_{H^{1}_{\overline{\mathbf{w}}}}\leq$ $\tilde{C}_{3}(\epsilon+\epsilon_{1})(\delta+\delta^{2})\sup_{s\in[0,\delta]}\|\overline{v}(t_{\diamond}+s)\|_{H^{1}_{\overline{\mathbf{w}}}}$
$+\tilde{C}_{3}\epsilon(\delta+\delta^{2})\sup_{s\in[0,\delta]}\big{|}f\big{(}\gamma(t_{\diamond}+s)\big{)}\big{|}.$
@@ 5 Long-time control
    We have $p\in[0,\frac{1}{4})$ together with $\epsilon>0$ and $\gamma>0$. In addition, we have $\delta=\epsilon^{-p}$ together with the inequalities
@@
    Given $T>0$, the function ${v}\in C([0,T],H^{2})\cap C^{1}([0,T],H^{-1})$ satisfies (2.7) while $\alpha\in C^{1}([0,T],{\mathbb{R}})$ solves (2.9). In addition, we have the inclusion
@@
Assume S1, S2 and the choice (5.1) for S3. If C3 and C4 are satisfied, and C5 holds for some $T>0$, then the bound
@@
Our main result here provides exponential control for the perturbation $v$ measured with respect to the time-varying spatial weight-functions $\mathbf{w}(t)$ defined in (5.3). It requires a-priori control over the terms in the expression
@@
which we have introduced for notational convenience. We note in particular that the condition (5.4) is satisfied if $K_{\epsilon,\gamma}(t)$ is sufficiently small.
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist constants $\delta_{8},C_{8}>0$ so that the following holds true. If C3 and C4 are satisfied, and C5 holds for some $T>0$, then the bound
@@
The proof of Proposition 5.2 is based on the evolution equation (2.7), which we rewrite as
@@
$M(t,v;\alpha,\Omega)=$ $\alpha^{-3}N({v})+2\frac{{\alpha}_{t}}{{\alpha}}\big{(}(1+\tfrac{1}{2}x\partial_{x})\phi_{c_{0}}+{v}\big{)}+\frac{{\Omega}_{t}}{{\alpha}}\partial_{x}(\phi_{c_{0}}+{v})$
$+\epsilon{f}(\gamma t)(\phi_{c_{0}}+{v}).$
@@
If furthermore $\delta>0$ satisfies $t+\delta\leq T$ and
@@
Writing $v(t)=v$, the estimate (5.8) follows by computing
@@
$\leq$ $2c_{0}\alpha^{-3}_{\mathrm{min}}C_{6}\delta\big{(}\|f\|_{\infty}\epsilon(1+w_{\mathrm{min}}^{-1/2}\epsilon_{1})+\epsilon_{1}^{2}\big{)}.$
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist a constant $C_{10}>0$ so that the following holds true. If C3 and C4 are satisfied, and C5 holds for some $T>0$, then for each $t\in[0,T]$ the bound
@@
Here we have used that $\int_{s}^{t}\alpha^{-3}(t^{\prime}){\mathrm{d}}t^{\prime}\geq\alpha^{-3}_{\mathrm{max}}(t-s)$. Using Lemma 3.2 and Lemma A.2, we estimate
@@
We furthermore use Lemma 4.3 to estimate
@@
$\|M(s,v(s);\alpha,\Omega)\|_{L^{2}_{\mathbf{w}(s)}}\leq$ $C_{9}\big{(}\epsilon(1+\epsilon_{1})+\epsilon_{1}+\epsilon_{2}+\epsilon_{1}^{2}\big{)}\|{v}(s)\|_{H^{1}_{\mathbf{w}(s)}}$
$+C_{9}\epsilon e^{-\gamma s}$
@@
Applying Lemma 5.4, we hence find
@@
For $\delta_{8}$ small enough, we have $\tilde{C}_{1}K_{\epsilon,\gamma}(T)<1/2$ and hence
$\sup_{0\leq t\leq T}e^{\gamma t}\|{v}(t)\|_{H^{1}_{\mathbf{w}(t)}}\leq 2\tilde{C}_{1}\big{(}\|\overline{v}_{*}\|_{H^{1}}+\|\overline{v}_{*}\|_{H^{1}_{w}}+\epsilon+\epsilon^{1-2p}\big{)}.$
@@ 6 Energy evolution
We set out to control the $L^{2}$-norm of the perturbation $\overline{v}$ by estimating the time-derivative of (6.1). The key point is that in the right-hand side of (6.2) below, the term $\|\overline{v}(t)\|_{L^{2}}^{2}$ comes with an integrable factor $\epsilon|f(\gamma t)|$, and the remaining terms are all integrable and small. Invoking Grönwall’s inequality this can subsequently be used to establish the desired control. We recall that $\delta_{6}$ is the constant introduced in Lemma 4.3.
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist a constant $C_{11}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C5 holds for some $T>0$ and $t\in[0,T]$, then the bound
@@
Let us write $u=u(t)$, $v=v(t)$ and $\overline{v}=\overline{v}(t)$ for brevity. In the presence of forcing via $\epsilon f(\gamma t)$, we observe that
@@
Combining (6.3) and (6.4), and estimating $|\alpha_{t}(t)+\tfrac{2}{3}\alpha(t)\epsilon f(\gamma t)|$ as in (4.6) leads to the estimate
$|\partial_{t}\|\overline{v}\|^{2}_{L^{2}}|=$ $|\partial_{t}\|u\|^{2}_{L^{2}}-\partial_{t}\|\phi_{c(t)}\|^{2}_{L^{2}}|$
$\leq$ $\tilde{C}_{1}\Big{(}\epsilon|f(\gamma t)|\|\overline{v}\|_{L^{2}}^{2}+\epsilon|f(\gamma t)|\|v\|_{L^{2}_{w_{\infty}}}+\|v\|^{2}_{L^{2}_{w_{\infty}}}\Big{)},$
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist a constant $C_{12}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C4 is satisfied, C5 holds for some $T>0$, and $\epsilon/\gamma\leq E_{\mathrm{max}}$, then the bound
@@
$\|\overline{v}\|^{2}_{L^{2}}\leq\Big{(}\|\overline{v}_{0}\|^{2}_{L^{2}}+\frac{C_{11}}{2\gamma}(\epsilon\epsilon_{1}+\epsilon_{1}^{2})\Big{)}e^{C_{11}\epsilon/\gamma}.$
Control in $H^{1}$ is established is established using the well-known fact that the soliton $\phi_{c}$ is a critical point of the functional $\mathcal{E}_{c}[u]=\tfrac{1}{2}c\|u\|_{L^{2}}^{2}+\mathcal{H}[u]$. This is reflected in the equality
@@
$\mathcal{J}(t)=$ $\mathcal{E}_{c(t)}[\phi_{c(t)}+\overline{v}(t)]-\mathcal{E}_{c(t)}[\phi_{c(t)}]$ (6.5)
$=$ $\tfrac{1}{2}c(t)\|\overline{v}(t)\|_{L^{2}}^{2}+\tfrac{1}{2}\|\partial_{x}\overline{v}(t)\|_{L^{2}}^{2}+\int_{{\mathbb{R}}}-\phi_{c(t)}\overline{v}^{2}(t)+\tfrac{1}{3}\overline{v}^{3}(t){\mathrm{d}}x,$
which generalises (6.1) in the sense that the right-hand side is $O(\overline{v}^{2})$. In fact, we note that $\|\overline{v}(t)\|_{H^{1}}^{2}$ can be bounded in terms of $|\mathcal{J}(t)|$ and vice versa.
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist a constant $C_{13}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C5 holds for some $T>0$ and $t\in[0,T]$, then
@@
in (6.5) yields
$|\mathcal{J}(t)|\leq$ $\tfrac{1}{2}\max\{c_{0}\alpha_{\mathrm{min}}^{-2},1\}\|\overline{v}(t)\|_{H^{1}}^{2}+\alpha^{-5}_{\mathrm{min}}\|e^{-2w_{\infty}x}\phi_{c_{0}}\|_{L^{\infty}}\|{v}(t)\|_{L_{w_{\infty}}^{2}}^{2}+\tfrac{\sqrt{2}}{3}\|\overline{v}(t)\|_{H^{1}}\|\overline{v}(t)\|_{L^{2}}^{2},$
which yields (6.7). On the other hand, rewriting (6.5) as
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist a constant $C_{14}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C5 hold for some $T>0$ and $t\in[0,T]$, then the bounds
@@
We once more abbreviate $u=u(t)$, $v=v(t)$ and $\overline{v}=\overline{v}(t)$, and compute that
@@
Rescaling the terms that contain the soliton, and introducing an exponential weight yields
@@
$\big{|}\partial_{t}\mathcal{H}[u]-\partial_{t}\mathcal{H}[\phi_{c(t)}]\big{|}\leq$ $(1+\sqrt{2})\epsilon|f(\gamma t)|\|\overline{v}\|_{H^{1}}^{2}+\tilde{C}_{3}\epsilon|f(\gamma t)|\big{(}\|v\|_{L^{2}_{w_{\infty}}}+\|v\|_{L^{2}_{w_{\infty}}}^{2}\big{)}+\tilde{C}_{3}\|v\|_{L^{2}_{w_{\infty}}}^{2}$
@@
Assuming S1, S2 and the choice (5.1) for S3, there exist a constant $C_{15}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C4 is satisfied, C5 holds for some $T>0$ and $\epsilon/\gamma\leq E_{\mathrm{max}}$, then the bounds
@@
which using Lemma 6.4 leads to the bound
@@
The result now follows by controlling $\epsilon_{2}$ via Lemma 6.2, controlling $|\mathcal{J}(0)|$ via (6.7) and applying item 1 of Lemma 2.2. ∎

### v2

@@ 4 Short-time control
$\left\|\int_{0}^{r}e^{(r-s)\mathcal{L}_{c(t_{\diamond})}}Q_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)\ {\mathrm{d}}s\right\|_{H^{1}_{\overline{\mathbf{w}}}}$
(3.3)${{\leq}}M\int_{0}^{\delta}e^{-\beta(\delta-s)}(\delta-s)^{-1/2}\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}{\mathrm{d}}s.$
@@
Corollary 4.4 can be used to derive that
@@
Furthermore, using $\int_{0}^{\delta}(\delta-s)^{-1/2}{\mathrm{d}}s=2\sqrt{\delta}$, we find that
@@
$P_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)=$ $\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t_{\diamond})}^{1}\big{\rangle}\partial_{x}\phi_{c(t_{\diamond})}$
$+\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t)}^{2}\big{\rangle}\partial_{c}\phi_{c(t_{\diamond})}$
where $\mathcal{L}_{c(t_{\diamond})}\partial_{x}\phi_{c(t_{\diamond})}=0$ and $\mathcal{L}_{c(t_{\diamond})}\partial_{c}\phi_{c(t_{\diamond})}=\partial_{x}\phi_{c(t_{\diamond})}$. Thus,
@@
$+\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t_{\diamond})}^{2}\big{\rangle}\partial_{c}\phi_{c(t_{\diamond})}$
$+\big{\langle}Y(t_{\diamond},s,\overline{v};c,\Omega),\eta_{c(t_{\diamond})}^{2}\big{\rangle}(\delta-s)\partial_{x}\phi_{c(t_{\diamond})},$
@@
$\|e^{(r-s)\mathcal{L}_{c(t_{\diamond})}}P_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{H^{1}_{\overline{\mathbf{w}}}}$
$\leq\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}\|\eta_{c(t_{\diamond})}^{1}\|_{L^{2}_{-{\overline{\mathbf{w}}}}}\|\partial_{x}\phi_{c(t_{\diamond})}\|_{H^{1}_{\overline{\mathbf{w}}}}$
$\quad+\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}\|\eta_{c(t_{\diamond})}^{2}\|_{L^{2}_{-{\overline{\mathbf{w}}}}}\|\partial_{c}\phi_{c(t_{\diamond})}\|_{H^{1}_{\overline{\mathbf{w}}}}$
$\quad+\|Y(t_{\diamond},s,\overline{v};c,\Omega)\|_{L^{2}_{\overline{\mathbf{w}}}}\|\eta_{c(t_{\diamond})}^{2}\|_{L^{2}_{-{\overline{\mathbf{w}}}}}(r-s)\|\partial_{x}\phi_{c(t_{\diamond})}\|_{H^{1}_{\overline{\mathbf{w}}}}$
@@
$\left\|\int_{0}^{r}e^{(r-s)\mathcal{L}_{c(t_{\diamond})}}P_{c(t_{\diamond})}Y(t_{\diamond},s,\overline{v};c,\Omega)\ {\mathrm{d}}s\right\|_{H^{1}_{\overline{\mathbf{w}}}}$
$\leq\tilde{C}_{3}(\epsilon+\epsilon_{1})(\delta+\delta^{2})\sup_{s\in[0,\delta]}\|\overline{v}(t_{\diamond}+s)\|_{H^{1}_{\overline{\mathbf{w}}}}+\tilde{C}_{3}\epsilon(\delta+\delta^{2})\sup_{s\in[0,\delta]}\big{|}f\big{(}\gamma(t_{\diamond}+s)\big{)}\big{|}.$
@@ 5 Long-time control
    We have $p\in[0,\frac{1}{4})$ together with $\epsilon\in(0,1]$ and $\gamma>0$. In addition, we have $\delta=\epsilon^{-p}$ together with the inequalities
@@
    Given $T>0$, the function ${v}\in C([0,T],H^{2})\cap C^{1}([0,T],H^{-1})$ satisfies (2.8), while $\alpha\in C^{1}([0,T],{\mathbb{R}})$ solves (2.10). In addition, we have the inclusion
@@
Assume S1, S2, and the choice (5.1) for S3. If C3 and C4 are satisfied, and C5 holds for some $T>0$, then the bound
@@
Our main result here provides exponential control for the perturbation $v$ measured with respect to the time-varying spatial weight-functions $\mathbf{w}(t)$ defined in (5.3). It requires a priori control [^2] This control is established in the proof of Theorem 1.1. over the terms in the expression
@@
which we have introduced for notational convenience. We note in particular that condition (5.4) is satisfied if $K_{\epsilon,\gamma}(t)$ is sufficiently small.
@@
Assuming S1, S2, and the choice (5.1) for S3, there exist constants $\delta_{8},C_{8}>0$ so that the following holds true. If C3 and C4 are satisfied, and C5 holds for some $T>0$, then the bound
@@
The proof of Proposition 5.2 is based on the evolution equation (2.8), in which we isolate the precarious term $\frac{{\alpha}_{t}}{{\alpha}}x\partial_{x}v$ as
@@
$M(t,v;\alpha,\Omega)\!=\!\alpha^{-3}N({v})\!+\!\frac{{\alpha}_{t}}{{\alpha}}\big{(}(2\!+\!x\partial_{x})\phi_{c_{0}}\!+\!2{v}\big{)}\!+\!\frac{{\Omega}_{t}}{{\alpha}}\partial_{x}(\phi_{c_{0}}\!+\!{v})\!+\!\epsilon{f}(\gamma t)(\phi_{c_{0}}+{v}).$
@@
If, furthermore, $\delta>0$ satisfies $t+\delta\leq T$ and
@@
Writing $v(t)=v$, estimate (5.8) follows by computing
@@
$\leq$ $2c_{0}\alpha^{-3}_{\mathrm{min}}C_{6}\delta\big{(}\|f\|_{\infty}\epsilon(1+w_{\mathrm{min}}^{-1/2}\epsilon_{1})+\epsilon_{1}^{2}\big{)}.\qed$
@@
Assuming S1, S2, and the choice (5.1) for S3, there exists a constant $C_{10}>0$ so that the following holds true. If C3 and C4 are satisfied, and C5 holds for some $T>0$, then for each $t\in[0,T]$, the bound
@@
Here, we have used that $\int_{s}^{t}\!\!\alpha^{-3}(t^{\prime}){\mathrm{d}}t^{\prime}\!\geq\!\alpha^{-3}_{\mathrm{max}}(t\!-\!s)$. Using Lemma 3.2 and Lemma A.2, we estimate
@@
Furthermore, we use Lemma 4.3 to estimate
@@
Recalling that $\delta=\epsilon^{-p}$ per C3, we have thus shown that
$e^{\gamma t}I_{1}\leq\tilde{C}_{3}K_{\epsilon,\gamma}(t)\sup_{s\in[0,t]}e^{\gamma s}\|{v}(s)\|_{H^{1}_{\mathbf{w}(s)}}.$
@@
$\|M(s,v(s);\alpha,\Omega)\|_{L^{2}_{\mathbf{w}(s)}}\leq C_{9}\big{(}\epsilon(1+\epsilon_{1})+\epsilon_{1}+\epsilon_{2}+\epsilon_{1}^{2}\big{)}\|{v}(s)\|_{H^{1}_{\mathbf{w}(s)}}+C_{9}\epsilon e^{-\gamma s}$
@@
Applying Lemma 5.4, we find
@@
For $\delta_{8}$ small enough, we have $\tilde{C}_{1}K_{\epsilon,\gamma}(T)<1/2$, and hence
$\sup_{0\leq t\leq T}e^{\gamma t}\|{v}(t)\|_{H^{1}_{\mathbf{w}(t)}}\leq 2\tilde{C}_{1}\big{(}\|\overline{v}_{*}\|_{H^{1}}+\|\overline{v}_{*}\|_{H^{1}_{w}}+\epsilon+\epsilon^{1-2p}\big{)}.\qed$
@@ 6 Energy evolution
We set out to control the $L^{2}$-norm of the perturbation $\overline{v}$ by estimating the time-derivative of (6.1). The key point is that, in the right-hand side of (6.2) below, the term $\|\overline{v}(t)\|_{L^{2}}^{2}$ comes with an integrable factor $\epsilon|f(\gamma t)|$, and the remaining terms are all integrable and small. Invoking Grönwall’s inequality, this can subsequently be used to establish the desired control. We recall that $\delta_{6}$ is the constant introduced in Lemma 4.3.
@@
Assuming S1, S2, and the choice (5.1) for S3, there exists a constant $C_{11}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C5 holds for some $T>0$ and $t\in[0,T]$, then the bound
@@
Let us write $u=u(t)$, $v=v(t)$, and $\overline{v}=\overline{v}(t)$ for brevity. In the presence of forcing via $\epsilon f(\gamma t)$, we observe that
@@
Combining (6.3) and (6.4) and estimating $|\alpha_{t}(t)+\tfrac{2}{3}\alpha(t)\epsilon f(\gamma t)|$ via (4.6) leads to the estimate
$|\partial_{t}\|\overline{v}\|^{2}_{L^{2}}|\!=\!|\partial_{t}\|u\|^{2}_{L^{2}}\!-\!\partial_{t}\|\phi_{c(t)}\|^{2}_{L^{2}}|\!\leq\!\tilde{C}_{1}\Big{(}\epsilon|f(\gamma t)|\|\overline{v}\|_{L^{2}}^{2}\!+\!\epsilon|f(\gamma t)|\|v\|_{L^{2}_{w_{\infty}}}\!+\!\|v\|^{2}_{L^{2}_{w_{\infty}}}\Big{)},$
@@
Assuming S1, S2, and the choice (5.1) for S3, there exists a constant $C_{12}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C4 is satisfied, C5 holds for some $T>0$, and $\epsilon/\gamma\leq E_{\mathrm{max}}$, then the bound
@@
$\|\overline{v}\|^{2}_{L^{2}}\leq\Big{(}\|\overline{v}_{0}\|^{2}_{L^{2}}+\frac{C_{11}}{2\gamma}(\epsilon\epsilon_{1}+\epsilon_{1}^{2})\Big{)}e^{C_{11}\epsilon/\gamma}.\qed$
Control in $H^{1}$ is established using the well-known fact that the soliton $\phi_{c}$ is a critical point of the functional $\mathcal{E}_{c}[u]=\tfrac{1}{2}c\|u\|_{L^{2}}^{2}+\mathcal{H}[u]$. This is reflected in the equality
@@
$\mathcal{J}(t)=$ $\mathcal{E}_{c(t)}[\phi_{c(t)}+\overline{v}(t)]-\mathcal{E}_{c(t)}[\phi_{c(t)}]$
$=$ $\tfrac{1}{2}c(t)\|\overline{v}(t)\|_{L^{2}}^{2}+\tfrac{1}{2}\|\partial_{x}\overline{v}(t)\|_{L^{2}}^{2}+\int_{{\mathbb{R}}}-\phi_{c(t)}\overline{v}^{2}(t)+\tfrac{1}{3}\overline{v}^{3}(t){\mathrm{d}}x,$ (6.5)
which generalizes (6.1) in the sense that the right-hand side is $O(\overline{v}^{2})$. In fact, we note that $\|\overline{v}(t)\|_{H^{1}}^{2}$ can be bounded in terms of $|\mathcal{J}(t)|$ and vice versa.
@@
Assuming S1, S2, and the choice (5.1) for S3, there exists a constant $C_{13}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C5 holds for some $T>0$ and $t\in[0,T]$, then
@@
in (6) yields
$|\mathcal{J}(t)|\leq$ $\tfrac{1}{2}\max\{c_{0}\alpha_{\mathrm{min}}^{-2},1\}\|\overline{v}(t)\|_{H^{1}}^{2}+\alpha^{-5}_{\mathrm{min}}\|e^{-2w_{\infty}x}\phi_{c_{0}}\|_{L^{\infty}}\|{v}(t)\|_{L_{w_{\infty}}^{2}}^{2}$
$+\tfrac{\sqrt{2}}{3}\|\overline{v}(t)\|_{H^{1}}\|\overline{v}(t)\|_{L^{2}}^{2},$
which yields (6.7). On the other hand, rewriting (6) as
@@
Assuming S1, S2, and the choice (5.1) for S3, there exists a constant $C_{14}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C5 holds for some $T>0$ and $t\in[0,T]$, then the bounds
@@
We once more abbreviate $u=u(t)$, $v=v(t)$, and $\overline{v}=\overline{v}(t)$, and compute that
@@
Rescaling the terms that contain the soliton and introducing an exponential weight yields
@@
$\big{|}\partial_{t}\mathcal{H}[u]-\partial_{t}\mathcal{H}[\phi_{c(t)}]\big{|}\leq$ $(1+\sqrt{2})\epsilon|f(\gamma t)|\|\overline{v}\|_{H^{1}}^{2}+\tilde{C}_{3}\epsilon|f(\gamma t)|\big{(}\|v\|_{L^{2}_{w_{\infty}}}+\|v\|_{L^{2}_{w_{\infty}}}^{2}\big{)}$
$+\tilde{C}_{3}\|v\|_{L^{2}_{w_{\infty}}}^{2}$
@@
Assuming S1, S2, and the choice (5.1) for S3, there exists a constant $C_{15}>0$ so that the following holds true for each $\epsilon,\gamma>0$. If C4 is satisfied, C5 holds for some $T>0$, and $\epsilon/\gamma\leq E_{\mathrm{max}}$, then the bounds
@@
which, using Lemma 6.4, leads to the bound
@@
The result now follows by controlling $\epsilon_{2}$ via Lemma 6.2, controlling $|\mathcal{J}(0)|$ via (6.7), and applying item 1 of Lemma 2.2. ∎